kenny carried quarters. doug carried nine dimes. together they had $2.65.how many quarters did kerry carry?
Answer by NancyLam(35)

let y=quater
0.1(9)+0.25y=$2.65
$0.9 + 0.25y=$2.65
y=$2.65-0.9/0.25
y=$1.75/0.25
y=$7
Check:0.1(9)+0.25(7)=$0.9+$1.75
= $2.65
the answer is: 7 quater
I hope it help